OpenStudy (anonymous):

f(x)=x^2+4x+1 and g(x)=3x+d... for what values of d does the equation f(x)=g(x) produce one solution?

OpenStudy (anonymous):

\[x^2+4x+1=3x+d\]subtract and get \[x^2+x+1-d=0\] it will have one solution if the discriminant \(b^2-4ac\) is 0. set \[1^2-4\times (1-d)=0\] and solve for \(d\)

OpenStudy (anonymous):

then what does it mean by "find the value of d when eq f(x)=g(x) produce 2 solutions"?

OpenStudy (anonymous):

two solutions means the discriminant is positive solve \[1-4(1-d)>0\] which should be real easy if you have already solved \[1-4(1-d)=0\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw